OpenStudy (anonymous):

Can someone check my answers please? Algebra.? 1. What is the degree of the monomial? 4x6 (1 point) 4 10 6** 24 2. What is the degree of the monomial? –2 (1 point) –2** –3 0 –1 3. What is the sum or difference? 5x8 + 9x8 (1 point) 45x8 –4x8** 14x16 14x8 4. What is the sum or difference? 5y4 – 9y4 (1 point) –45y4 14y4 –4y4 –4y8** 5. Write the polynomial in standard form. Then name the polynomial based on its degree and number of terms. 5x + 2x4 – 4 (1 point) 5x – 4 + 2x4; not a polynomial** 2x4 + 5x – 4; fourth-degree

OpenStudy (anonymous):

5x + 2x4 – 4; quadratic monomial 2x4 + 5x – 4; cubic binomial 6. Simplify the sum. (4u3 + 5u2 + 6) + (8u3 – 7u + 2) (1 point) 12u3 + 5u2 – 7u + 8** 12u3 – 2u2 – 7u + 8 –4u3 – 7u2 + 5u – 8 –4u3 + 5u2 – 7u + 8 7. Simplify the sum. (5u3 + 5u2 + 3) + (2u3 – 3u + 8) (1 point) 3u3 – 3u2 + 5u – 11 3u3 + 5u2 – 3u + 11** 7u3 + 2u2 – 3u + 11 7u3 + 5u2 – 3u + 11 8. Find the GCF of the terms of the polynomial. 32x2 + 28x6 (1 point) 4x2 x2 4x6** 28x2 9. Find the GCF of the terms of the polynomial. 16x4 + 38x3 (1 point) 16x** x3 2x4 2x3 10. Simplify the product using the distributive property. (2h + 2)(2h + 6) (1 point) 4h2 – 8h – 12** 4h2 + 8h – 12 4h2 + 16h + 12 4h2 – 16h + 12 11. Simplify the product using FOIL. (3x – 5)(3x + 7) (1 point) 9x2 + 6x – 35** 9x2 + 36x + 35 9x2 – 6x – 35 9x2 – 36x + 35 12. What is a simpler form of the expression? (3k + 4)(2k2 – 3k – 5) (1 point) 6k3 + 17k2 – 3k – 20 6k3 – k2 – 27k – 20 6k3 – k2 + 3k – 20** 6k3 + 27k2 – k – 20 15. What is a simpler form of the following expression? (3m2 – 6)(3m2 + 6) (1 point) 9m3 – 36 9m2 – 36** 9m4 + 36 9m4 – 36 16. What is the factored form of the expression? d2 + 5d + 4 (1 point) (d – 1)(d + 4) (d + 1)(d + 4)** (d – 1)(d – 4) (d + 1)(d – 4) 17. What is the factored form of the expression? x2 + xy – 2y2 (1 point) (x + y)(x – 2y) (x + y)(x + 2y) (x – y)(x – 2y) (x – y)(x + 2y)** 18. What is the factored form of the expression? 56x2 + 85x + 24 (1 point) (8x + 3)(7x – 8)** (8x + 3)(7x + 8) (8x – 3)(7x – 8) (8x – 3)(7x + 8) 19. What is the factored form of the expression? 6x2 + 17x + 12 (1 point) (3x – 4)(2x + 3) (3x – 4)(2x – 3) (3x + 4)(2x + 3)** (3x + 4)(2x – 3) 20. What is the factored form of the expression? 6g2 + 5g – 6 (1 point) (3g – 2)(2g – 3)** (3g + 2)(2g – 3) (3g + 2)(2g + 3) (3g – 2)(2g + 3) 21. What is the factored form of the expression? 28g2 – 37gh – 21h2 (1 point) (4g + 7h)(7g – 3h) (4g + 7)(7g + 3h2) (4g – 7)(7g + 3) (4g – 7h)(7g + 3h)** 22. What is the factored form of the expression? 168y2 – 99y – 42 (1 point) 3(8y – 7)(7y + 2) (8y – 7)(21y + 6)** (8y + 7)(7y – 2) 3(8y + 7)(7y + 2) 23. Find the radius of a circle with an area of CE_pi_large(16x2 + 24x + 9). Use the formula A = CE_pi_larger2 to find the area of a circle. (1 point) 3x – 4 9x – 16** 16x + 9 4x + 3 24. What is the factored form of the expression? 80x2 + 120x + 45 (1 point) 5(4x + 3)2 5(16x + 9)2 5(3x + 3)2** 5(3x + 9)2 25. A refrigerator has a volume given by the trinomial y3 + 3y2 – 10y. What are the possible dimensions of the refrigerator? Use factoring. (1 point) y – 2, –y – 2, and –y – 5 y – 5, y – 5, and y + 5** y, y + 2, and y – 5 y, y – 2, and y + 5

OpenStudy (anonymous):

Can someone please check these? They need to be done by tomorrow and I don't want a bad grade.

OpenStudy (unklerhaukus):

1✓,2✓ try 3,4,5 again 6✓, try 7,8,9,10,11,12,15 again 16✓,17✓, try 18,19,20, again 21✓,22✓ i dont understand 23 try 24,25 again

OpenStudy (unklerhaukus):

\[ax^n+bx^n=(a+b)x^n\]\[(a+b)(c+d)=ac+ad+bc+bd\]

OpenStudy (anonymous):

Thank you! c: